### v1

# BiasFreeBench: a Benchmark for Mitigating Bias in Large Language Model Responses

Abstract
Existing studies on bias mitigation methods for large language models (LLMs) use diverse baselines and metrics to evaluate debiasing performance, leading to inconsistent comparisons among them. Moreover, their evaluations are mostly based on the comparison between LLMs’ probabilities of biased and unbiased contexts, which ignores the gap between such evaluations and real-world use cases where users interact with LLMs by reading model responses and expect fair and safe outputs rather than LLMs’ probabilities. To enable consistent evaluation across debiasing methods and bridge this gap, we introduce BiasFreeBench, an empirical benchmark that comprehensively compares eight mainstream bias mitigation techniques (covering four prompting-based and four training-based methods) on two test scenarios (multi-choice QA and open-ended multi-turn QA) by reorganizing existing datasets into a unified query-response setting. We further introduce a response-level metric, Bias-Free Score, to measure the extent to which LLM responses are fair, safe, and anti-stereotypical. Debiasing performances are systematically compared and analyzed across key dimensions: the prompting vs. training paradigm, model size, and generalization of different training strategies to unseen bias types. We release our benchmark, aiming to establish a unified testbed for bias mitigation research.
Warning: This paper contains content that may be offensive and upsetting.
1 Introduction
Table 1: Existing Debiasing Evaluation.
Debiasing Techniques
Have both training-
and prompting-
based baselines?
Evaluate bias
in LLM response?
DAMA (Limisiewicz et al., 2024)
Furniturewala et al. (2024)
BiasDPO (Allam, 2024)
FAST (Chen et al., 2025)
BiasEdit (Xu et al., 2025)
FairSteer (Li et al., 2025)
Self-Debiasing (Gallegos et al., 2025)
Modern large language models, such as ChatGPT (OpenAI, 2023), display biased behaviors when interacting with humans, despite being trained to align with human values through reinforcement learning from human feedback (Goldfarb-Tarrant et al., 2023; Gallegos et al., 2024; Oba et al., 2024; Naous et al., 2024; Echterhoff et al., 2024). Recent debiasing techniques for modern LLMs have been proposed, but they adopt inconsistent evaluation setups as shown in Table 1. On the one hand, varying and inconsistent baselines are chosen, making results difficult to compare. On the other hand, most evaluations are based on LLM internal probability by comparing the likelihoods of biased and unbiased contexts (Nangia et al., 2020; Nadeem et al., 2021), rather than directly assessing bias in the model’s generated responses. This creates a gap between research practice and real-world usage, where LLMs are used in a query–response setting and users directly interpret their outputs. To date, however, there has been little systematic and consistent comparison of various bias mitigation techniques for LLM responses.
To address these limitations, we introduce BiasFreeBench, a new benchmark that provides a comprehensive and empirical comparison of debiasing methods for modern LLMs by directly evaluating bias in their responses, as shown in Figure 1. BiasFreeBench is constructed with three unique designs:
Figure 1: BiasFreeBench  comprehensively compare prompting-based and training-based techniques to mitigate bias in LLM responses. They are evaluated on QA-based bias datasets with a response-level metric, Bias-Free Score.
  1. 1.
    Multi-dimensional comparisons of bias mitigation techniques: BiasFreeBench includes eight representative debiasing techniques spanning prompting-based (Self-Reflection, Self-Awareness, Self-Help, Chain-of-Thought) and training-based (DPO, SFT, Safe Alignment, Task Vector) strategies, enabling comprehensive analysis across methods and settings. In this study, seven LLMs with different sizes, including instruction-tuned LLMs, reasoning LLMs, and commercial LLMs, are investigated. Debiasing performances are analyzed under the implementation paradigms, model sizes, and bias types.
  2. 2.
    Unified test scenarios tailored for modern LLMs: BiasFreeBench reformats existing bias evaluation datasets into the query-response style. For example, we adapt BBQ (Parrish et al., 2022), a multiple-choice QA benchmark with gold bias annotations, into the single-turn query-response format to reflect real-world LLM usage. It also incorporates FairMT-Bench, a multi-turn conversational QA dataset with open-ended questions without ground truths, which also supports evaluation under both short and long-context dialogue settings.
  3. 3.
    A new response-level metric design: To better capture bias in LLM outputs for aligning with human needs in practical use, we propose the Bias-Free Score, a novel metric that directly assesses bias in model outputs by quantifying the proportion of responses that are safe, fair, and anti-stereotypical.
We evaluate these techniques along three axes: $1)$ the effectiveness of prompting- vs. training-based techniques, $2)$ performance scaling with model size, and $3)$ the generalization across different bias types. Our empirical findings show that prompting-based methods are consistently more effective than training-based methods. A simple prompt intervention, such as Self-Awareness, can effectively reduce response bias and show consistent improvements with larger model sizes. Meanwhile, some training techniques like DPO exhibit strong generalization across bias types, suggesting that training on a single bias category can yield broader fairness benefits. We present BiasFreeBenchas a unified testbed for rigorous and fair evaluation of bias mitigation methods, and hope our findings provide practical insights to guide future research on response-level debiasing in LLMs.
2 Related Work
Previous debiasing techniques for relatively small languages, like BERT (Devlin et al., 2019) and GPT2 (Radford et al., 2019) have various forms. Some approaches fine-tune models using counterfactual data that swap identity terms (Zmigrod et al., 2019; Lu et al., 2020; Xu et al., 2022) while others modify internal representations by projecting them onto unbiased subspaces (Liang et al., 2020; Shi et al., 2024). To improve efficiency, alternative efficient debiasing fine-tuning strategies are proposed (Gira et al., 2022). Biased prompts and prompting techniques (Gehman et al., 2020; Sheng et al., 2020; Guo et al., 2022) are introduced to help models adjust their biases. On the one hand, some methods based on representation projection (Liang et al., 2020; Ravfogel et al., 2020) remove bias representations from models but do not fundamentally alter their internal biases without modifying model parameters. On the other hand, Kumar et al. (2023); Yu et al. (2023); Chen et al. (2025); Xu et al. (2025) try to use adapters and machine unlearning or editing to debias models parametrically. They are mainly evaluated on and designed for likelihood-based text modeling (Meade et al., 2022). For example, two stereotype datasets, StereoSet (Nadeem et al., 2021) and Crows-Pairs (Nangia et al., 2020), with bias annotation measure debiasing performance based on the likelihood of bias attribute terms or whole sentences with bias attributes.
Works about debiasing evaluation and bias mitigation for modern chat LLMs have emerged recently. CEB (Wang et al., 2025), BiasAlert (Fan et al., 2024), and BiasGuard (Fan et al., 2025b) investigate fairness evaluation for LLM responses. Echterhoff et al. (2024); Oba et al. (2024); Furniturewala et al. (2024); Kamruzzaman & Kim (2024); Gallegos et al. (2025) design different prompting strategies to directly control LLMs’ outputs. Walsh et al. (2025); Borah & Mihalcea (2024) trys multi-agent collaboration. Model editing (Ilharco et al., 2023; Limisiewicz et al., 2024), unlearning (Dige et al., 2024) and steering methods (Li et al., 2025) for debiasing LLMs are explored. Post-training, including DPO (Allam, 2024; Zhang et al., 2025), SFT (Raza et al., 2024), and Safe RLHF (Dai et al., 2024) is also used to guide LLMs to output safe information. However, to date, there is no comprehensive comparison among them. Therefore, we construct a benchmark, BiasFreeBench, to study them comprehensively.
3 Techniques in BiasFreeBench
3.1 Evaluation Task: Mitigating Bias in LLM Response
Given a query $q\in\mathcal{Q}$, a large language model $f_{\theta}:\mathcal{Q}\rightarrow\mathcal{R}$ will output a response $r\in\mathcal{R}$. In a single-turn dialogue, there is one query $q$ and one response $r$. In a multi-turn chat, there are $N$ sequential query-response pairs, i.e. $\{(q_{1},r_{1}),(q_{2},r_{2}),...,(q_{N},r_{N})\},q_{i}\in\mathcal{Q},r\in\mathcal{R},i\in[1,N]$. Each response $r_{i}$ is generated based on previous dialogue history $\{(q_{1},r_{1}),...,(q_{i-1},r_{i-1}),q_{i}\}$. Modern LLMs can give biased responses (Wang et al., 2025; Fan et al., 2025a). For example, when we ask Llama-3.1-8B-Instruct (Dubey et al., 2024) the following question (a sample from BBQ (Parrish et al., 2022)),
The office is shared between a person who jogs after work and a person who uses a wheelchair. Who is productive? ‘The person who jogs after work’ or ‘The person who uses a wheelchair’? Please choose one answer.
it gives a biased response ‘The person who jogs after work’. Concentrating on the harmful bias in LLM’s interactive responses $r$, we construct a comprehensive benchmark BiasFreeBench to evaluate existing bias mitigation techniques, including mainstream prompting-based and training-based methods.
3.2 Prompting-based Mitigation
In our benchmark, we include four zero-shot prompting-based bias mitigation techniques (see examples with dialogue histories in Appendix D.4), as shown in Figure 2.
Self-Awareness
Inspired by the way that humans use to reduce cognitive bias, Echterhoff et al. (2024) adds hints and warnings for a certain type of bias to a query, which makes LLMs aware of the bias type when responding to the query. Specifically, there is a Self-Awareness instruction at the end of the query, where Bias Type refers to a particular bias type, such as gender, age, disability, etc., or a general bias type, e.g., ‘social’ and ‘cultural’, so that this method can flexibly adapt to any bias type.
Figure 2: Instructions for the prompting-based debiasing methods
Self-Reflection
Reflection (Shinn et al., 2023) is an effective method to improve reasoning performance for an LLM agent. Similar to reflection, self-debiasing (Gallegos et al., 2025) reprompts a LLM to remove bias from its first response and give a new one. According to them, we design a Self-Reflection debiasing instruction. After an LLM gives a response to a query, we ask the LLM with this instruction to reflect on (i.e., recheck) the response and remove potentially ‘recognized’ bias by giving a response again, which helps the LLM to be aware of bias and maintain consistency with the query and initial response.
Self-Help
The inputs to an LLM sometimes contain biased information, directly leading to a biased response. Therefore, it is important to mitigate bias in a query. Besides using instructions to control LLM responses, we also investigate LLMs’ potential to discover and remove bias in prompts, including both system prompts and input queries. Following Echterhoff et al. (2024), we ask an LLM to rewrite prompts to avoid giving biased responses. Then, the rewritten system prompt and query will be used to query the LLM in a new session. This self-help mechanism enables LLMs to autonomously refine potentially biased inputs, reducing the reliance on human intervention, but two forward passes are necessary.
CoT
Chain-of-Thought (CoT) has been demonstrated as an effective method for enhancing LLM reasoning capabilities (Wei et al., 2022). Following Ganguli et al. (2023), we instruct the model to engage in step-by-step thinking for avoiding biased responses.
Figure 3: Four training-based bias mitigation techniques explored in BiasFreeBench.
3.3 Training-based Mitigation
BiasFreeBench  explored two kinds of training-based bias mitigation techniques: $i)$ alignment-based training, including DPO, SFT, and safe alignment and ii)ii) post-editing method such as Task Vector, as shown in Figure 3.
SFT
Supervised Fine-Tuning (SFT) tunes LLMs on curated datasets to align with desired behaviors, such as instruction following. We leverage SFT to directly refine the biased patterns in model responses by learning about bias-free responses. Given a query ‘The physicist was getting ready to go to work.’, for example, a LLM is directly trained to give an anti-stereotypical response ‘She is a young woman with perfect vision.’ (a gender-profession stereotype from Nadeem et al. (2021)).
DPO
As an RLHF-based training method to align LLMs with human preferences, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is leveraged in this work to tune LLMs to generate bias-free responses while discouraging biased outputs. For instance, given a query q=q=The physicist was getting ready to go to work.’, DPO guides an LLM to favor generating an anti-stereotypical response $r_{w}=$She is a young woman with perfect vision.’ and give a penalty for a stereotypical response $r_{l}=$He is an older white male with glasses.’.
Safe Alignment
Safe alignment trains LLMs to align with ethical and safety principles and prevent harmful, biased, or inappropriate outputs. Specifically, we use two phases of Safe RLHF (Dai et al., 2024). In the first phase, a reward model (RM) and a cost model (CM) are trained on a helpfulness dataset and a harmlessness dataset, respectively. In the second safe reinforcement learning phase, the RM and CM estimate the value of human preference for helpfulness and harmlessness, respectively, and a modern LLM is trained based on these two values to align with safe human values.
Task Vector
Task Vector (Ilharco et al., 2023) is a model editing method used to mitigate biases learned during previous training. Firstly, an LLM $\theta_{pre}$ is trained via SFT to output a biased response given a query, which will obtain a biased LLM $\theta_{biased}$. Secondly, a bias vector $\tau$ is calculated as the element-wise difference between the weights of $\theta_{biased}$ and $\theta_{pre}$, i.e., $\tau=\theta_{biased}-\theta_{pre}$. Finally, it updates the LLM $\theta_{pre}$ in the opposite direction of $\tau$, i.e., $\theta_{biasfree}=\theta_{pre}-\tau$ to remove the bias effect introduced by the bias vector and obtain a bias-free model $\theta_{biasfree}$.
4 Implementation Design
4.1 Model and Training Setups
In this study, we investigate seven LLMs, including $i)$ instruction-tuned LLMs: Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Qwen-2.5-7B-Instruct (Yang et al., 2024), and deepseek-llm-7b-chat (Bi et al., 2024), ii)ii) reasoning LLMs: DeepSeek-R1-Distill-Llama-8B (Guo et al., 2025) and Qwen3-8B (Yang et al., 2025), $iii)$ commercial LLM: gpt-4o-mini [^1] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ They are debiased with four prompting-based techniques, and four training-based techniques (§3.1) and evaluated on two bias evaluation datasets (§4.2). We use the intersentence portion of StereoSet (Nadeem et al., 2021) as the training data for SFT, DPO, and Task Vector. Specifically, each training sample consists of a context as a query $q$, a stereotypical response $r_{l}$, and an anti-stereotypical response $r_{w}$. In DPO, we use ($q$, $r_{l}$, $r_{w}$) as a sample where $r_{w}$ is the positive output and $r_{l}$ is the negative output following Dige et al. (2024). In SFT and Task Vector, we use ($q$, $r_{w}$). Safe Alignment pipeline is implemented with Safe RLHF (Dai et al., 2024). More details are in Appendix B.
4.2 Evaluation Datasets and Metrics
We evaluate the effectiveness of bias mitigation techniques in two dataset settings under a unified query-response framework, which aligns with real-world human-LLM interaction: (1) single-turn QA with gold bias annotations, e.g., BBQ (Parrish et al., 2022), and (2) multi-turn conversational QA where LLMs generate open-ended responses, e.g., FairMT-Bench (Fan et al., 2025a). A new metric Bias-Free Score (BFS) is also proposed to measure response-level bias in for LLMs. The detailed evaluations for each setting are elaborated as follow:
BBQ
is a bias benchmark for multi-choice QA. Each sample consists of a context, a question, and three candidate answers with gold bias annotation: $i)$ biased responses, ii)ii) anti-stereotypical responses, and $iii)$ UNKNOWN where a LLM gives a safe response, such as ’it cannot be determined without enough information’, ’I cannot give an answer because the query is harmful’, ’I choose both $i)$ and ii)ii).’. The context can provide ambiguous or disambiguous information required to answer the question. Following Dige et al. (2024); Gallegos et al. (2025) and Wang et al. (2025), we only use samples with ambiguous contexts to evaluate potential biases. More details are described in Appendix C. To form a conversational prompt for each sample, we concatenate the context, question, option $i)$, ii)ii), and an instruction at the end to enforce LLMs to follow the query. Examples are shown in §3.1 and Appendix D.4. Based on our bias mitigation objective that LLMs are expected to give bias-free responses, we define BFS of BBQ as:
$BFS_{\text{BBQ}}=\frac{N_{ii)}+N_{iii)}}{N_{i)}+N_{ii)}+N_{iii)}}$ (1)
where $N_{\text{annotation}}$ is the number of samples that belong to the annotation.
FairMT-Bench
is a fairness benchmark designed for open-domain, multi-turn dialogue scenarios (details in Appendix D.4). Each sample contains a user query from five consecutive turns of dialogue. For each turn, the LLM generates an open-text response. Following the original setup as Fan et al. (2025a), we evaluate bias mitigation performance by measuring bias in the LLM’s response of the last turn. Examples are shown in Appendix D.4. Since no gold annotations are available, we classify LLM responses for FairMT-Bench into two categories: $i)$ biased, ii)ii) UNKNOWN (such as fair and safe responses, cannot answer without enough information, and refusal to answer a potentially harmful query). Similarly, the Bias-Free Score of FairMT-Bench is calculated as
$BFS_{\text{FairMT-Bench}}=\frac{N_{ii)}}{N_{i)}+N_{ii)}}$ (2)
4.3 Evaluation with LLMs and Tools
For each query in BBQ and FairMT, we first obtain the response from the LLM. Then we use GPT-4o mini [^2] https://platform.openai.com/docs/models/gpt-4o-mini, Llama-Guard-3-8B [^3] https://huggingface.co/meta-llama/Llama-Guard-3-8B, and Moderation API [^4] https://platform.openai.com/docs/guides/moderation as judgers. For BBQ, we prompt GPT-4o-mini three times to identify which gold bias annotation each response aligns best with and aggregate the results via majority vote (Wang et al., 2023). For FairMT-Bench, we classify responses using GPT-4o-mini ($i)$ biased vs. ii)ii)unknown), Llama-Guard (ii)ii) safe vs. $i)$ unsafe), and Moderation API ($i)$ toxic vs. ii)ii) non-toxic), and again apply majority voting to obtain the final label. The judgment prompts of GPT-4o-mini for each dataset are shown in Appendix B.2. We also conduct human checks (see Appendix B.3), where we find that the LLM judgment is very available since it achieves 100% agreement with humans for BBQ (Cohen’s kappa (McHugh, 2012) = 1.0), and 94% agreement with humans for FairMT-Bench (Cohen’s kappa = 0.7).
5 Experimental Results
5.1 Main Discussion on Debiasing Techniques
The results of debiasing performance are shown in Table 2 and 3.
[…]
Llama-3.1 Mistral Qwen2.5 dp-llm-chat dp-R1-Llama Qwen3 gpt-4o-mini
Vanilla 52.41 81.24 44.28 53.94 46.75 50.25 46.86
Prompting
Self-Awareness 52.55 91.60 46.69 73.72 57.34 61.31 56.54
Self-Reflection 82.66 90.79 58.36 70.10 80.91 91.31 79.20
Self-Help 95.52 92.09 80.69 85.48 71.91 78.44 92.23
CoT 82.82 92.63 87.24 61.94 96.11 91.98 92.48
Average (Prompting) 78.39 91.78 68.25 72.81 76.57 80.76 80.11
Training
SFT 52.11 81.17 44.40 46.32 43.84 40.27 -
DPO 58.56 85.86 43.41 60.77 53.54 45.90 -
Task Vector 82.77 89.95 64.56 93.88 49.61 47.31 -
Safe RLHF 46.09 47.30 38.75 44.82 - -
Average (Training) 59.88 76.07 47.78 61.45 49.00 44.49 -
[…]
Llama3.1 Mistral Qwen2.5 dp-llm-chat dp-R1-Llama Qwen3 gpt-4o-mini
Vanilla 76.84 73.30 58.83 66.61 77.80 79.90 66.33
Prompting
Self-Awareness 89.20 92.73 94.24 89.37 90.70 95.92 93.61
Self-Reflection 82.96 90.64 84.09 88.36 95.13 96.86 95.58
Self-Help 78.83 86.85 66.67 72.87 74.72 82.56 71.73
CoT 94.40 95.93 95.18 94.72 98.56 98.56 97.89
Average (Prompting) 86.35 91.54 85.05 86.33 89.78 93.48 89.70
Training
SFT 82.10 78.74 65.73 68.45 71.71 81.85 -
DPO 82.54 82.14 59.63 71.22 85.69 83.33 -
Task Vector 80.61 86.12 63.82 67.26 60.11 83.98 -
Safe RLHF 88.74 40.11 44.44 64.83 - - -
Average (Training) 83.50 71.78 58.41 67.94 72.50 83.05 -
5.1.1 Analysis: Prompting-based Mitigations
From Table 2 and 3, we noticed that CoT achieves the best debiasing performance (i.e., the highest BFSes) in most cases on both BBQ and FairMT-Bench, indicating that exposing (potentially biased) reasoning helps mitigate biased responses. In contrast, other prompting-based methods yield more varied performance. Comparing the BFS(%) improvement with Self-Help on BBQ (up to 43.11) and FairMT-Bench (up to 7.84), we observe that Self-Help performs strongly in the BBQ-like setting where the context is short and has the hint of the options, but its effectiveness drops significantly on very long contexts of FairMT-Bench because rewriting coherent and benign prompts becomes more challenging as the context length increases (Liu et al., 2024). For instance, as shown in Figure 16 and 17, a rewritten query can change the meaning of the original query, leading to an unrelated response (3.81% responses semantically misaligned with the original queries). Instead, Self-Awareness yields the second-best performance on FairMT-Bench in most cases, with less computation cost (Appendix D.2) as it does not require a second pass of querying LLM as Self-reflection and self-help, which illustrates that Self-Awareness offer both solid performance and greater efficiency.
5.1.2 Analysis: Training-based Mitigations
By comparing the alignment training methods in Table 2 and 3, we notice $1)$ DPO yields better debiasing performance than SFT in most cases maybe because SFT learns from safe-only examples, leading the model to mimic safe responses, while DPO learns the preference by comparing safe and unsafe behaviors, leading to better discrimination and generalization. $2)$ Although Safe Alignment adds an explicit constraint on harmfulness, it often leads to large BFS drops over two datasets. The conjecture is that the helpfulness reward in Safe RLHF tends to make the LLM decisive, inhibiting ambiguous responses (fewer UNKNOWN responses are observed, shown in Appendix D.3), indicating the challenges of finding a nuanced balance between helpfulness and harmfulness using constrained optimization. $3)$ The post-editing method, Task Vector, achieves better debiasing than alignment methods. However, we found that it also sacrifices the general performance after editing the model, as shown in the next paragraph.
General Capabilities Retention.
We investigate whether training-based debiasing methods will harm the general capabilities of LLMs. We evaluate the understanding, reasoning, and truthfulness abilities of LLMs on three benchmark datasets, BoolQ (Clark et al., 2019), COPA (Gordon et al., 2012), and TruthfulQA (Lin et al., 2022), respectively, using OpenCompass [^5] https://github.com/open-compass/opencompass, and report the accuracy difference between the vanilla LLM and the debiased one in Table 4. The results show tiny performance differences for DPO, SFT, and Safe RLHF. However, Task Vector decreases LLM general capabilities, indicating the challenge of editing models without overly changing them, as also noticed by other model editing methods (Gu et al., 2024; Gupta et al., 2024).
Table 4: Accuracy changes for general capabilities. BoolQ and COPA: Accuracy (%). TruthfulQA: BLEU Accuracy.
Vanilla SFT DPO Task Vector Safe RLHF Vanilla SFT DPO Task Vector Safe RLHF
Llama-3.1-8B-Instruct Mistral-7B-Instruct-v0.3
BoolQ 85.38 -0.03 +0.34 -22.57 -1.95 81.99 0.00 -0.55 -10.99 +0.85
COPA 94.00 0.00 -1.00 -34.00 +3.00 95.00 0.00 0.00 -34.00 +1.00
TruthfulQA 0.29 0.00 +0.01 -0.11 0.00 0.29 0.00 0.00 -0.20 -0.01
Qwen2.5-7B-Instruct deepseek-llm-7b-chat
BoolQ 85.11 +0.03 +0.30 -14.53 +2.11 82.14 -0.46 -0.61 -11.65 +0.92
COPA 93.00 +1.00 +1.00 -13.00 0.00 94.00 -2.00 -2.00 -15.00 -1.00
TruthfulQA 0.31 0.00 0.00 -0.06 -0.03 0.29 -0.02 -0.01 -0.13 +0.01
5.1.3 Comparison: Prompting vs. Training
By comparing the average BFS of prompting-based and training-based techniques, we notice that among the eight debiasing techniques we explored, prompting-based bias mitigation techniques generally demonstrate stronger performance compared to training-based methods. Many studies (Chen et al., 2022; Xie et al., 2024; Xu et al., 2024; Cheng et al., 2024) have shown that when presented with conflicting information, LLMs prioritize the contextual input over their internal parametric knowledge. This aligns with the test case of debiasing, where in all prompting-based methods, the input prompts introduce bias-free (anti-stereotypical) cues that are contrastive to the model’s internal stereotypical knowledge. Therefore, the prompts effectively override biases embedded in its parametric knowledge. In contrast, training-based methods attempt to generally modify the model’s internal representations, which is challenging because biases are parametrically scattered in model weights, even deeply ingrained in only a few modules (Limisiewicz et al., 2024; Xu et al., 2025; Chen et al., 2025) and difficult to fully erase without affecting the general knowledge stored in model weights.
5.2 Debiasing with Different Model Sizes
Figure 4: Mean and standard deviation of BFS (%) across 4 prompting-based and 3 training-based methods on different sizes of Qwen2.5.
To investigate the effectiveness of different bias mitigation techniques across various model sizes of LLMs, we evaluate 4 prompting-based (Self-Awareness, Self-Reflection, Self-Help, and CoT) and 3 training-based techniques (SFT, DPO, and Task Vector) on 5 different sizes of Qwen2.5. We draw the average performance line in each category and use shades to show the variance in Figure 4. We observe that prompting-based bias mitigation techniques generally outperform training-based techniques across different model sizes, but with greater variance than training-based techniques, as the shaded areas indicate. What’s more, as model size increases, the BFS of prompting-based methods steadily improves, suggesting that larger models are better at using prompt engineering to reduce bias. In contrast, training-based methods maintain relatively stable performance across model sizes. The conjecture is that the effectiveness of prompting benefits from the greater knowledge and reasoning capacity of larger models, while training-based approaches rely more on the quality and coverage of the training data than on model scale.
5.3 Training with Different Bias Type
(a) Various bias types in different LLMs.
[…]
6 Conclusion
Previous works for mitigating LLMs’ bias use diverse and inconsistent evaluation patterns. Therefore, in this work, we construct BiasFreeBench, an empirical benchmark to systematically evaluate bias mitigation techniques through LLM responses. We assess four prompting-based and four training-based debiasing techniques across two QA-style bias datasets using seven types of LLMs of varying sizes. To align with real-world human usage, we focus on bias in LLM responses. A new response-level metric, Bias-Free Score, is introduced to provide a direct measurement of mitigation effectiveness in LLM outputs. We hope that this benchmark can serve as a unified testbed for bias mitigation methods, and our findings can inspire further research in designing more effective bias mitigation systems for LLM responses.
Ethics Statement
We recognize the potential societal impacts of our work that debiasing techniques in BiasFreeBench can be used immorally to make LLMs give biased responses, which is harmful to society. We advocate for the responsible use of our method in ways that benefit the whole society and minimize harm.
Reproducibility statement
The open-source LLM weights we used are from HuggingFace (Wolf et al., 2019). The tools and commercial model are used with their official code and API. All detailed experimental settings are provided in Section 3, 4, 5, and the Appendix to ensure reproducibility. There is no private information in our research. All data and codes for the experiments are publicly available in https://github.com/xxupiano/BiasFreeBench.
Appendix A LLM Usage
In this work, LLMs were only used for the following two purposes:
  1. 1.
    Experiments: implementations of debiasing, evaluations, llm-as-a-judge, as shown in Section 4 and 5.
  2. 2.
    Helping paper writing: grammar check and writing polishing.
Appendix B Experimental Details
B.1 Implementation
Inference
Except for the second session in Self-Help, the system prompt of all prompting-based methods and evaluations for training-based methods on BBQ is ‘You are a helpful assistant.’. As for FairMT-Bench, we follow the official system prompt, which is ‘You are a helpful AI assistant skilled in multi-round conversations. Please answer the user’s questions based on the context of the discussions, ensuring coherence and relevance. Answer in less than **128** words.’ except for Self-Help. The system prompts for Self-Help are shown in Figure 13, 16, and 17. Since there are no bias type annotations in FairMT-Bench, Bias Type is replaced as ‘social’ / ‘socially’ in the instructions among different prompting strategies. All inference is implemented by vllm (Kwon et al., 2023) with two 48G RTX A6000.
Training Data
We use the inter-sentence part of StereoSet as the training data following Dige et al. (2024). The reason is that this is the only suitable bias dataset with DPO-form input-output pairs and rich meta-information. Though a new bias dataset for DPO, BiasDPO (Allam, 2024), is constructed, it has only about 1,000 samples without meta-information. However, StereoSet is in a long-tailed distribution. To investigate whether the unbalanced data will influence the debiasing performance, we first adopt a weighted sampling strategy to balance the training data. Specifically, we calculate the inverse frequency of each bias type and assign higher sampling probabilities to underrepresented categories, which ensures that each bias type is adequately represented in the sampled dataset and mitigates the effects of data imbalance while maintaining the overall dataset size. Detailed numbers of them are shown in Table 5. Then, both the long-tailed data and the balanced data are used to implement SFT, and DPO. The results in Table 6 show that training with the balanced dataset outperforms training with the unbalanced dataset in 62.5% of the cases. Therefore, SFT, DPO, and Task Vector in this work were implemented with the balanced training data except the analysis experiments in §5.3. In §5.3, the training with a single bias type of data is conducted with the original unbalanced data from StereoSet, while the training with mixed bias types of data is conducted with the balanced data.
Table 5: Distribution of different bias types in the original StereoSet and our balanced training data.
Bias Type # Origin # Balanced
Race 3,923 2,129
Gender 993 2,141
Profession 3,262 2,100
Religion 319 2,127
Total 8,497 8,497
Table 6: Bias-Free Score (%) of balanced training data vs original unbalanced training data.
Llama-3.1 Mistral Qwen2.5 deepseek-llm
BBQ
Unbalanced SFT 50.68 45.76 41.68 46.95
Balanced SFT 52.11 41.17 44.40 46.32
Unbalanced DPO 55.71 85.81 42.63 58.36
Balanced DPO 58.56 85.86 43.41 60.77
FairMT-Bench
Unbalanced SFT 80.85 71.38 68.41 69.55
Balanced SFT 82.10 78.74 65.73 68.45
Unbalanced DPO 85.88 80.08 60.07 70.59
Balanced DPO 82.54 82.14 59.63 71.22
Training
We implement DPO with LoRA (Hu et al., 2022), SFT with LoRA, Task Vector training with full SFT by LLaMA-Factory (Zheng et al., 2024). We didn’t implement the SFT phase of the original Safe RLHF (Dai et al., 2024) since modern chat LLMs had already undergone instruction tuning. In the RL phase of Safe RLHF, we only use PKU-SafeRLHF-10K [^7] https://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-10K for GPU time limitations. Two 48GB NVIDIA RTX A6000 are utilized for DPO and SFT Eight 80G NVIDIA H100 are used for Task Vector and Safe RLHF.
FairMT-Bench Distribution
FairMT-Bench evaluates the fairness of LLM across three stages: context understanding, interaction fairness, and instruction trade-offs. The detailed statistics are presented in Table 7.
Table 7: Statistics of FairMT-Bench across three stages.
Stage Task Num.
Context Understanding Scattered Questions 1,889
Anaphora Ellipsis 2,152
Interaction Fairness Jailbreak Tips 864
Interference Misinformation 1,655
Instruction Trade-offs Fixed Format 1,980
Negative Feedback 1,655
Total 10,195
Table 8: SFT and DPO Settings for LLaMA-Factory (More details are shown in the code.)
SFT DPO
Hyper-parameter Value Hyper-parameter Value
GPU 2 * RTX A6000 GPU 2 * RTX A6000
Training Batch Size per GPU 16 Training Batch Size per GPU 16
Gradient Accumulation Steps 4 Gradient Accumulation Steps 4
Learning Rate 8.0e-6 Learning Rate 8.0e-6
Train Epochs 20 Train Epochs 20
LR Scheduler Type cosine LR Scheduler Type cosine
Warmup Ratio 0.1 Warmup Ratio 0.1
bf16 TRUE bf16 TRUE
Load Best Model at End TRUE Load Best Model at End TRUE
B.2 LLM for Judgment
Prompts used for judgment are in Figure 6, Figure 8, and Figure 8. To mitigate positional bias in reference selection, we randomly shuffle the order of references in each calling. As for BBQ, with the prompt in Figure 6 and majority vote, only about 0.12% samples have the LLM responses that don’t semantically align with the queries so we discard them when calculating BFS. As for FairMT-Bench, after evaluation with the prompt in Figure 8, we observe that there are about 4% LLM responses that don’t semantically align with the queries, especially for Self-Help. We conjecture that the reason is that it’s difficult for current LLMs to deal with very long contexts. Therefore, we use the prompt in Figure 8 to call gpt-4o-mini twice and then apply majority vote to determine if the LLM response of one sample semantically align with the query. Same as BBQ, we discard these samples when calculating BFS.
Figure 6: Judgment Prompt for BBQ.
Figure 7: Judgment Prompt for FairMT-Bench.
Figure 8: Judgment Prompt for correlation between a query and an LLM response.
B.3 Manual Checks for LLM Judgment
Four NLP graduate students as reviewers manually checked the LLM judgments. For BBQ, each reviewer examined 100 samples, and they found that all the LLM judgments were correct. For FairMT-Bench, each reviewer examined 100 samples. They were required to categorize each LLM response into one of the following categories: $i)$ biased/stereotypical; ii)ii) unbiased/fair/safe/cannot answer/not enough information; $iii)$ does not semantically answer the query / does not follow the query / is not related to the query. Then we calculate the judgment agreement as the percentage of samples whose response annotation from the LLM/Tool judgment is the same as the human judgment. Meanwhile, we calculate Cohen’s Kappa between gpt-4o-mini and humans.
Appendix C BBQ Settings
C.1 Bias Type Distribution
Table 9: Distribution of ambiguous samples in BBQ.
Bias Type Num. Bias Type Nums.
Disability Status 778 Age 1,840
Gender Identity 2,828 Nationality 1,540
Physical Appearance 788 Race / Ethnicity 3,440
Socioeconomic Status (SES) 3,432 Religion 600
Sexual Orientation 432 Total 15,678
C.2 Gold Bias Annotation
According to the answers that contribute to a positive bias score from BBQ paper (Parrish et al., 2022), we extract the gold bias annotation for ambiguous contexts as follows: $i)$ biased response: the bias target in the negative context and the non-target in the non-negative context; ii)ii) anti-stereotypical response: the non-target in the negative context and the bias target in the the non-negative context; $iii)$ UNKNOWN: the UNKNOWN answer option.
C.3 Only using ambiguous contexts
In the same way as Dige et al. (2024); Gallegos et al. (2025); Wang et al. (2025), we don’t consider disambiguous contexts. Without disambiguous information, a model will only rely on its stereotypical assumptions, whereas the detailed information for target groups in disambiguous contexts will make LLM pay attention to the factuality and distract from bias assumptions. Moreover, because all other bias datasets only have one kind of query for two target groups, except BBQ and we want to construct a benchmark that can be generalized to most bias datasets, we only explore the commonly-used data format in our paper.
C.4 The link between $BFS_{\text{BBQ}}$ and original BBQ metrics
According to BBQ paper, we keep using $n$ to represent the number of examples that fall into each response group, therefore $n_{\text{biased\_ans}}$ represents the number of model outputs that reflect the targeted social bias (i.e., the bias target in negative contexts and the non-target in non-negative contexts), $n_{\text{anti\_ans}}$ represents the number of model outputs that non-target anti-stereotype, and $n_{\text{non-unk\_outputs}}$ is the total number of model outputs that are not Unk (i.e., all target and non-target outputs). As define in BBQ, the Bias score in disambiguated contexts:
$s_{\textsc{Dis}}$ $=2\left(\frac{n_{\text{biased\_ans}}}{n_{\text{non-unk\_outputs}}}\right)-1$ (3)
$=\frac{2n_{\text{biased\_ans}}-(n_{\text{biased\_ans}}+n_{\text{anti\_ans}})}{n_{\text{non-unk\_outputs}}}$ (4)
$=\frac{n_{\text{biased\_ans}}-n_{\text{anti\_ans}}}{n_{\text{non-unk\_outputs}}}$ (5)
Therefore the Bias score in ambiguous contexts defined in BBQ paper is
$s_{\textsc{Amb}}$ $=(1-\text{accuracy})s_{\textsc{Dis}}$ (6)
$=\left(1-\frac{n_{\text{unk\_outputs}}}{n_{\text{non-unk\_outputs}}+n_{\text{unk\_outputs}}}\right)$ (7)
$\times\left(\frac{n_{\text{biased\_ans}}-n_{\text{anti\_ans}}}{n_{\text{non-unk\_outputs}}}\right)$ (8)
$=\left(\frac{n_{\text{non-unk\_outputs}}}{n_{\text{non-unk\_outputs}}+n_{\text{unk\_outputs}}}\right)$ (9)
$\times\left(\frac{n_{\text{biased\_ans}}-n_{\text{anti\_ans}}}{n_{\text{non-unk\_outputs}}}\right)$ (10)
$=\frac{n_{\text{biased\_ans}}-n_{\text{anti\_ans}}}{n_{\text{all\_outputs}}}$ (11)
A perfect model without any bias will output UNK for all ambiguous examples, resulting accuracy=1, therefore, the $s_{\textsc{Amb}}=0$; A extremely biased model will have no UNK output, resulting in accuracy=0, $s_{\textsc{Amb}}=1\times s_{\textsc{Dis}}=s_{\textsc{Dis}}$. Therefore, the score $s_{\textsc{Amb}}$ ranges in (1,1)(-1,1).
Let’s define the rescaled $s^{\prime}_{\textsc{Amb}}$ as:
$s^{\prime}_{\textsc{Amb}}$ $=\frac{1+s_{\textsc{Amb}}}{2}$ (12)
$=\frac{n_{\text{biased\_ans}}+0.5\times n_{\text{unk\_ans}}}{n_{\text{all\_outputs}}}$ (13)
whose ranges in $(0,1)$. Then “Bias-Free Score” in the original BBQ context on ambiguous examples is:
$BFS_{\text{BBQ\_ori}}$ $=1-s^{\prime}_{\textsc{Amb}}$ (14)
$=\frac{n_{\text{anti\_ans}}+0.5\times n_{\text{unk\_ans}}}{n_{\text{all\_outputs}}}$ (15)
whereas our BFS score is
$BFS_{\text{BBQ}}=\frac{n_{\text{anti\_ans}}+1\times n_{\text{unk\_ans}}}{n_{\text{all\_outputs}}}$ (16)
Therefore, our BFS can be viewed as a reweighted version of the bias-free score under the evaluation metrics of the original BBQ paper. Conceptually, it is also intuitive that in an ambiguous setup where no explicit context is provided to infer the answer, a neutral response (what we call Unk here) should be equally preferred as selecting anti-stereotypical options, instead of being less preferred in the $BFS_{\text{BBQ\_ori}}$ (as it weights $n_{\text{unk\_ans}}$ by 0.5).
According to the formulations of $BFS_{\text{BBQ}}$ in C.4, we find that our intuitive metric $BFS_{\text{BBQ}}$ is the reweighted version of $BFS_{\text{BBQ\_ori}}$. To explore whether this reweighting will affect the robustness of experimental results, we report the “Bias-Free Score” in the original BBQ context $BFS_{\text{BBQ\_ori}}$ in Table 10. We observe that our key conclusions remain consistent across both versions, though our proposed $BFS_{\text{BBQ}}$ (with weight = 1.0) highlights the trends more clearly.
Table 10: $BFS_{\text{BBQ\_ori}}$ (%)
Llama3.1 Mistral Qwen2.5 dp-llm-chat dp-r1-llama Qwen3 gpt-4o-mini
Vanilla 40.09 46.17 38.82 45.03 39.62 37.98 37.05
Prompting
Self-Awareness 41.51 48.62 41.68 47.21 41.01 42.87 40.58
Self-Reflection 49.56 48.41 44.26 48.31 48.07 49.64 57.43
Self-Help 48.94 47.83 46.53 48.52 45.23 44.75 48.34
CoT 47.68 48.99 48.44 45.44 49.29 48.49 48.27
Average 46.92 48.46 45.23 47.37 45.90 46.44 48.66
Training
SFT 39.81 46.18 38.95 45.96 38.34 39.45 -
DPO 39.94 46.75 38.25 46.45 41.55 37.23 -
Task Vector 46.59 48.03 43.15 49.13 39.31 38.29
Safe RLHF 37.94 47.12 38.27 44.41 - - -
Average 41.07 47.02 39.66 46.49 39.73 38.32 -
(a) Bias-Free Score (%) of bias mitigation techniques with BBQ.
(b) Bias-Free Score (%) of bias mitigation techniques FairMT-Bench.
Figure 9: Bias-Free Scores Across Model Sizes.
Appendix D More Experimental Results
D.1 Model Size
The BFS of different bias mitigation techniques among Qwen2.5 with different sizes are shown in Figure 9.
D.2 Token Costs
Table 11 shows. comparisons of token costs, which is from evaluating Llama-3.1-8B-Instruct on BBQ (2 48G RTX A6000 with vllm, inference batch size=8) with the multi-pass prompt methods (Self-Reflection and Self-Help), and single-pass methods (CoT and Self-Awareness).
Table 11: Comparison of # token usage across prompting techniques.
1-round Input 1-round Output 2-round Input 2-round Output Sum
Single-pass
Self-Awareness 1,094,775 110,727 1,205,502
CoT 1,251,555 956,870 2,208,425
Multi-pass
Self-Reflection 905,413 111,867 291,896 274,793 1,583,969
Self-Help 2,748,395 992,971 857,739 2,302,045 6,901,150
D.3 UNKNOWN Rate
Table 12 reports the ratio of samples with UNKNOWN responses on BBQ. We observe that SFT and Safe RLHF lead to very low ratios of UNKNOWN responses. As for SFT, training only with anti-stereotypical will lead an LLM to be prone to give fewer UNKNOWN responses. As for Safe RLHF, the trade-off between helpful and UNKNOWN responses is a challenging problem.
Table 12: The ratio of samples with UNKNOWN responses on BBQ (%). S-Aware: Self-Awareness. S-Refl: Self-Reflection. S-Help: Self-Help. TV: Task Vector.
Vanilla S-Aware S-Refl S-Help CoT SFT DPO TV Safe RLHF
Llama-3.1-8B-Instruct 24.64 22.08 66.21 93.15 70.27 12.76 37.25 72.36 16.30
Mistral-7B-Instruct-v0.3 70.14 85.96 84.76 88.52 87.29 0.38 78.21 83.84 0.38
Qwen2.5-7B-Instruct 10.93 10.03 28.20 68.31 77.61 10.89 10.31 42.82 0.95
deepseek-llm-7b-chat 17.83 53.02 43.59 73.93 33.00 0.91 28.64 89.50 0.82
D.4 Cases
Here are some real cases that include prompts and LLM responses for prompting-based methods: Figure 11, 11, 12, 13,14, 15, 16, 17.
Figure 10: Self-Awareness example on BBQ.
Figure 11: CoT example on BBQ.
Figure 12: Self-Reflection example on BBQ.
(a) Session 1
(b) Session 2
Figure 13: Self-Help example on BBQ.
Figure 14: Self-Awareness example on FairMT-Bench.
Figure 15: Self-Reflection example on FairMT-Bench.
Figure 16: Session 1 of a Self-Help example on FairMT-Bench.
Figure 17: Session 2 of a Self-Help example on FairMT-Bench.